# A Survey on Text-Driven 360-Degree Panorama Generation

### v2

@@ IV-B1 NAR-Based
An early attempt in this paradigm, ImmerseGAN [14], utilized a GAN-based inpainting architecture [79] for this task. To achieve text-guided outpainting, ImmerseGAN adopted a pre-trained discriminative network to produce a latent vector representing the given textual description. This latent vector subsequently guided the generator to produce a 360-degree panorama semantically consistent with the text prompt. More recent approaches have primarily focused on leveraging the power of pre-trained latent diffusion models (LDMs) for their strong image generation priors acquired from training on large-scale datasets.
@@
PanoDiff [42], the first LDM-based method for text-driven NFoV outpainting, was trained on the SUN360 [48] dataset. It initially converts the input NFoV images into partial panoramas with visibility masks, and then employs a ControlNet-based LDM [37] for text-guided panorama completion. To ensure geometric continuity at the borders of the generated panorama, PanoDiff further implements a circular padding scheme during inference. Similarly, Diffusion360 [58] adopts a ControlNet-based LDM [37] to generate 360-degree panoramas from perspective images and textual descriptions. However, instead of circular padding, Diffusion360 leverages a circular blending strategy during the denoising and VAE decoding stages for improved boundary continuity of the generated 360-degree panorama. Recognizing that a single network (as employed in PanoDiff and Diffusion360) often struggles to simultaneously learn the inherent 360-degree panoramic distortion and perform content completion, PanoDecouple [24] designs a decoupled diffusion model. This framework decouples the NFoV outpainting process into distortion guidance and content completion. Furthermore, it develops a distortion correction loss, leveraging the proposed Distort-CLIP, to explicitly constrain the distinct 360-degree panoramic distortion.
@@
Depart from the aforementioned methods [24, 42, 58], which predominantly process and generate 360-degree panoramas using an equirectangular representation throughout their networks, several recent studies have explored leveraging alternative representations for 360-degree panoramic synthesis. CubeDiff [44], inspired by multi-view diffusion models [62, 61], generates 360-degree panoramas in cubemap format. This cubemap representation enables CubeDiff to more effectively leverage the diffusion priors learned by the LDM from extensive perspective images during the generation process. CubeDiff fine-tunes a pre-trained LDM on a mixed dataset of Structured3D [47], Pano360 [19], Polyhaven [5], and Humus [6], using a single conditional view (NFoV image) and textual embeddings as input. To model inter-face relationships within the cubemap, CubeDiff inflates all 2D attention layers in the LDM, enabling effective learning of inter-view dependencies. Similarly, [23] produces six overlapping perspective images, each with a 110110 field of view, using a Transformer [75] architecture fine-tuned with LoRA on the SUN360 [48] and Structured3D [47] datasets. These overlapping images are subsequently stitched together to synthesize the final 360-degree panoramic image.
@@ IV-B2 AR-Based
AOG-Net [60] introduces an autoregressive framework, building upon a pre-trained LDM [63], to progressively outpaint NFoV images into complete panoramas under textual guidance. This stepwise approach enhances the generation of fine-grained visual content and improves alignment with the input textual descriptions. Moreover, AOG-Net incorporates a global-local conditioning mechanism to integrate text, omni-visual, NFoV and omni-geometry guidances at each autoregression step, thereby ensuring consistency across the generated panorama. AOG-Net is trained on Laval Indoor [45] and Laval Outdoor [46] for indoor and outdoor scenarios, respectively. Following the training dataset settings in AOG-Net, OPa-Ma [4] uses a pre-trained LDM with Mamba [3], a state-space model known for its efficiency in handling long sequences, to iteratively outpaint local regions in each step. It introduces two modules: the Visual-textual Consistency Refiner, which enhances input utilization during generation, and the Global-local Mamba Adapter, which ensures global coherence across the generated panorama. In contrast to methods leveraging pre-trained LDMs (such as AOG-Net and OPa-Ma), the recent PAR [25] adopts a text-conditioned masked autoregressive model [78] for 360-degree panoramic image outpainting. Guided by text descriptions, PAR incorporates its designed dual-space circular padding and a translation consistency loss to improve output quality.
IV-C Comparisons
To systematically evaluate the strengths and weaknesses of representative methods of different kinds, we conduct both qualitative and quantitative benchmarks on Text-Only Generation (Sec. IV-A) and Text-Driven NFoV Outpainting (Sec. IV-B) methods with publicly available inference codes and model checkpoints. For Text-Only Generation, we compare Text2Light [56], Diffusion360 [58], StitchDiffusion [57] and PanFusion [59]. For Text-Driven NFoV Outpainting, we compare PanoDiff [42] and Diffusion360.
@@ IV-C Comparisons
IV-C1 Qualitative Comparison
Fig. 6 presents a visual comparison of the four text-only generation methods. To facilitate the assessment of visual and semantic continuities at the image borders, we generate ”swapped” versions by exchanging the left and right halves of each generated image. Red dashed boxes are employed in these swapped images to highlight the boundary regions. Among the evaluated methods, Text2Light exhibits obvious seams in the highlighted areas of its swapped images, indicating a failure to maintain visual and semantic continuities at the borders. While Diffusion360 effectively mitigates such observable seams and preserves visual continuity, it occasionally demonstrates local semantic discontinuity at the borders, where content across image boundaries lacks consistent logical or meaningful coherence. In contrast, both StitchDiffusion and PanFusion achieve strong visual and semantic continuities at the borders of their synthesized images.
Fig. 7 displays the visual comparison for the two text-driven NFoV outpainting approaches. Similar to the text-only comparison, swapped image versions were created to scrutinize the visual and semantic continuities at the borders. Both PanoDiff and Diffusion360 demonstrate commendable visual and semantic continuities at the borders of the produced images.
IV-C2 Quantitative Comparison
@@
To quantitatively assess the performance of these representative methods, we conduct a comprehensive comparison using metrics outlined in Sec. III-C. To ensure an unbiased evaluation of the generalizability of the methods, we employ an out-of-domain dataset, ODI-SR [1], on which none of the models have been explicitly trained. For generating text descriptions, we utilize BLIP2 [18] to create textual captions for the 360-degree panoramas included in the ODI-SR dataset. These generated text prompts serve as inputs for both Text-Only Generation and Text-Driven NFoV Outpainting tasks. To simulate NFoV images, we first project the equirectangular 360-degree panoramas from the ODI-SR dataset into a cubemap format and then extract the front face of each cubemap. The original 360-degree panoramas from ODI-SR are designated as real images and used as ground truth for the computation of evaluation metrics.
@@
The quantitative results obtained from this comparative evaluation, across the seven evaluation metrics, are presented in Tab. IV. To provide insights into the computational efficiency of each method, we also report the inference time and GPU memory required to generate a 1024512 360-degree panorama on a consistent machine equipped with an RTX A6000 GPU. These quantitative findings effectively delineate the strengths and weaknesses of the evaluated approaches, offering valuable guidance for future research in this domain.
V Emerging 3D Applications
@@
Recent advances in text-driven 360-degree panorama generation [57, 58, 59] have catalyzed innovative methods for reconstructing 360-degree 3D scenes from textual descriptions. 360-degree panoramic images inherently capture both global contexts and geometric constraints of a scene, making them an essential intermediate representation for 3D scene generation. Consequently, recent text-driven 360-degree 3D scene generation methods [31, 30, 33, 32, 27] utilize 360-degree panorama generation to bridge the gap between text prompts and 360-degree 3D scene reconstruction.
@@
As depicted in Fig. 8, these methods typically use a two-stage process: (1) 360-Degree Panorama Generation: generating a 360-degree panorama from the input text prompt using a fine-tuned LDM [63], and (2) 3D Scene Reconstruction: inferring a 3D representation, typically with 3D Gaussian Splatting (3DGS) [26], from the generated panorama and corresponding multi-view perspective images. Tab. V provides a comparative summary.
@@
Within this framework, emerging methods are primarily differentiated by their choice of (a) 360-degree panorama generators and their (b) strategies for extracting and utilizing 3D information. For instance, FastScene [31] and HoloDreamer [30] both employ Diffusion360 [58] to generate the initial 360-degree panorama depicting a scene from a given text prompt. FastScene [31] then synthesizes multi-view panoramas of this scene for specific camera poses using Coarse View Synthesis and Progressive Novel View Inpainting. With these synthesized multi-view panoramas, FastScene introduces Multi-View Projection to get their perspective views. The point clouds derived from these views are then used as input for 3DGS to reconstruct the 3D scene. HoloDreamer [30] enhances the Diffusion360-generated panorama with two distinct ControlNet-based LDMs [37] and a super-resolution network to create a high-resolution, stylized output. Subsequently, HoloDreamer initializes 3D Gaussians using point clouds derived from a reverse equirectangular projection of the high-resolution panorama combined with its corresponding depth information. Finally, a two-stage 3DGS optimization process is developed to refine the scene rendering, resulting in the desired 3D scene reconstruction.
@@
Furthermore, certain methods deviate from the reliance on Diffusion360 [58]. DreamScene360 [33] utilizes StitchDiffusion [57] to generate multiple 360-degree panorama candidates and then employs a self-refinement process to select the optimal candidate for initializing panoramic 3D Gaussians with a 3D geometric field. To facilitate visual feature correspondences between different views and maintain geometric consistencies during the 3DGS optimization process, semantic and geometric regularizations are applied. In contrast, SceneDreamer360 [32] uses a fine-tuned PanFusion [59] generator, coupled with a super-resolution module from [58], to produce a high-resolution (6K) panorama aligned with the input text prompt. It then uses optimization-based viewpoint selection to extract multi-view images, which are subsequently used for improved point cloud initialization, ultimately leading to 3DGS-based scene reconstruction.
@@
Other methods explore alternative panorama generation techniques. LayerPano3D [27] begins by generating four orthogonal perspective views with a fine-tuned text-to-image model [21]. These initial views are then combined with text-guided inpainting [63], and further processed by using a fine-tuned Diffusion360 [58] model to outpaint the polar regions, resulting in a reference 360-degree panorama. To handle occlusions in complex scenes, LayerPano3D [27] decomposes the reference panorama into multiple depth-based layers and utilizes a fine-tuned inpainter [59] to complete unseen content at each layer. These inpainted, layered panoramas then provide supervision for panoramic 3D Gaussian scene optimization.

### v3

@@ IV-B1 NAR-Based
An early attempt in this paradigm, ImmerseGAN [14], uses a GAN-based inpainting architecture [79] for this task. To achieve text-guided outpainting, ImmerseGAN adopts a pre-trained discriminative network to produce a latent vector representing the given textual description. This latent vector subsequently guides the generator to produce a 360-degree panorama semantically consistent with the text prompt. More recent approaches have primarily focused on leveraging the power of pre-trained latent diffusion models (LDMs) for their strong image generation priors acquired from training on large-scale datasets.
@@
PanoDiff [42], the first LDM-based method for text-driven NFoV outpainting, is trained on the SUN360 [48] dataset. It initially converts the input NFoV images into partial panoramas with visibility masks, and then employs a ControlNet-based LDM [37] for text-guided panorama completion. To ensure geometric continuity at the borders of the generated panorama, PanoDiff further implements a circular padding scheme during inference. Similarly, Diffusion360 [58] adopts a ControlNet-based LDM [37] to generate 360-degree panoramas from perspective images and textual descriptions. However, instead of circular padding, Diffusion360 leverages a circular blending strategy during the denoising and VAE decoding stages for improved boundary continuity of the generated 360-degree panorama. Recognizing that a single network (as employed in PanoDiff and Diffusion360) often struggles to simultaneously learn the inherent 360-degree panoramic distortion and perform content completion, PanoDecouple [24] introduces a decoupled diffusion model as its core contribution. This framework separates the NFoV outpainting process into distortion guidance and content completion. While this modular design is effective, it increases model complexity by requiring a separately trained Distort-CLIP model. Contrastingly, SpND [82] incorporates structural prior information from 360-degree panoramic images processed through a spherical network into its diffusion model to guide the 360-degree panoramic image outpainting process.
@@
Depart from the aforementioned methods [24, 42, 58], which predominantly process and generate 360-degree panoramas using an equirectangular representation throughout their networks, several recent studies have explored leveraging alternative representations for 360-degree panoramic synthesis. CubeDiff [44], inspired by multi-view diffusion models [62, 61], generates 360-degree panoramas in cubemap format. This cubemap representation enables CubeDiff to more effectively leverage the diffusion priors learned by the LDM from extensive perspective images during the generation process. CubeDiff fine-tunes a pre-trained LDM on a mixed dataset of Structured3D [47], Pano360 [19], Polyhaven [5], and Humus [6], using a single conditional view (NFoV image) and textual embeddings as input. Its central innovation is the inflation of 2D attention layers in the LDM into 3D attention layers, enabling the model to explicitly learn inter-face dependencies. This effectively trades the ERP distortion problem for a complex inter-view consistency challenge. Similarly, [23] produces six overlapping perspective images, each with a 110110 field of view, using a Transformer [75] architecture fine-tuned with LoRA on the SUN360 [48] and Structured3D [47] datasets. These overlapping images are subsequently stitched together to synthesize the final 360-degree panoramic image. To address the computational redundancy and resolution constrains introduced by these overlapping images, DreamCube [83] designs a multi-plane synchronization approach, enabling seamless and consistent RGB-D cubemap generation without overlaps as its main contribution.
@@ IV-B2 AR-Based
AOG-Net [60] introduces an autoregressive framework, building upon a pre-trained LDM [63], to progressively outpaint NFoV images into complete panoramas under textual guidance. This stepwise approach enhances the generation of fine-grained visual content and improves alignment with the input textual descriptions. Its key innovation is a global-local conditioning mechanism that integrates multiple guidance signals at each step to ensure consistency across the generated panorama. AOG-Net is trained on Laval Indoor [45] and Laval Outdoor [46] for indoor and outdoor scenarios, respectively. Following the training dataset settings in AOG-Net, OPa-Ma [4] uses a pre-trained LDM with Mamba [3], a state-space model known for its efficiency in handling long sequences, to iteratively outpaint local regions in each step. It introduces two modules: the Visual-textual Consistency Refiner, which enhances input utilization during generation, and the Global-local Mamba Adapter, which ensures global coherence across the generated panorama. In contrast to methods leveraging pre-trained LDMs (such as AOG-Net and OPa-Ma), the recent PAR [25] adopts a text-conditioned masked autoregressive model [78] for 360-degree panoramic image outpainting. Guided by text descriptions, PAR incorporates its designed dual-space circular padding and a translation consistency loss to improve output quality. For all AR-based methods, this step-by-step generation enforces strong logical coherence but at the cost of slower inference and a susceptibility to error propagation.
IV-C Analysis of Fundamental Design Choices
The methods surveyed above make fundamentally different design choices that entail significant trade-offs in performance, efficiency, and quality. A critical examination of these choices is essential for understanding the current research landscape and identifying future directions. Below, we systematically analyze the trade-offs and implications of three core architectural decisions: (1) the choice of panoramic representation, (2) the generation framework, and (3) the strategy for adapting pre-trained models.
IV-C1 Panoramic Representation – Simplicity vs. Prior Compatibility
As detailed in Sec. III-A, methods typically adopt ERP, CMP, or MPP. This selection reflects a trade-off between representational simplicity and compatibility with the generative priors of large pre-trained text-to-image models.
Implications of Choosing ERP: ERP offers a single continuous image, making it directly compatible with standard 2D architectures such as U-Nets and Transformers. However, this choice presents two significant challenges. (1) Standard 2D convolutions are not inherently periodic, so to enforce seamlessness at the 360-degree wrap-around boundary, methods must incorporate special treatments such as circular padding (PanoDiff [42]), circular blending (Diffusion360 [58]), or dedicated stitching steps (StitchDiffusion [57]). (2) ERP introduces non-uniform texel density (geometric distortions), stretching content near the poles and complicating learning of accurate 360-degree panoramas. Some methods address this with spherical manifold convolution (SMGD [81]) or a distortion guidance branch (PanoDecouple [24]), albeit at additional computational cost.
Implications of Choosing CMP/MPP: The decision to use a collection of perspective views (CMP or MPP) aims to leverage the priors of pre-trained models more effectively, as these models were trained predominantly on low-distortion perspective images. Representing the 360-degree panoramic scene in this format, as done in CubeDiff [44] and [23], maximizes compatibility and often yields higher-fidelity details. The trade-off is a shift in complexity: the problem of distortion is replaced by the problem of inter-view consistency. This necessitates specialized mechanisms to ensure seamlessness between views, such as inflating 2D attention layers to model 3D relations (CubeDiff [44]) or employing complex multi-step stitching and inpainting pipelines (PanoFree [41]), which can increase computational overhead and introduce new types of boundary artifacts.
IV-C2 Generation Framework – NAR-based vs. AR-based
The underlying mechanism for generating 360-degree panoramic content represents a key architectural choice, centered on a trade-off between generation speed and global coherence.
Non-Autoregressive (NAR-based): Most recent methods are built upon non-autoregressive LDMs [63], in which the entire latent representation of the 360-degree panorama is denoised in parallel over a series of steps. The primary advantage of this approach is computational efficiency, enabling relatively fast generation of full panoramas, as demonstrated by Diffusion360 [58]. The main drawback is the difficulty of ensuring global coherence across the vast 360-degree scene. Because the whole image is generated in a single holistic process, the model may struggle to maintain consistent long-range context, sometimes producing mismatched regions or repetitive textures.
Autoregressive (AR-based): In contrast, an autoregressive framework generates the panorama sequentially, explicitly conditioning each new region on the content generated in previous steps. This step-by-step process naturally enforces both local and global consistency, and methods like AOG-Net [60] and PAR [25] often exhibit strong logical coherence and fine-grained detail. The trade-off for this improved consistency is a substantial increase in inference time, as the sequential process is inherently slower than parallel generation. Moreover, these methods can suffer from error propagation, where low-quality generation in an early step can negatively impact the quality of all subsequent parts.
IV-C3 Adaptation of Pre-trained Models – Fine-tuning vs. Training-Free
Since no large-scale generative model is natively trained on 360-degree panoramic image-text pairs, researchers have to adapt models originally trained on standard perspective image-text datasets. Three distinct strategies have emerged, spanning a spectrum of trade-offs among computational cost, data requirements, and model specialization.
Full/Substantial Fine-Tuning: One strategy is to fine-tune a large portion of the model parameters on a specialized 360-degree panoramic image-text dataset. For instance, Diffusion360 [58] leverages the DreamBooth [38] technique to adapt a pre-trained LDM [63]. This approach enables the model to learn the unique geometric properties and data distributions of 360-degree panoramas, often leading to strong performance. However, the trade-offs are significant: it is usually computationally expensive and time-consuming, requires a substantial panoramic dataset, and may induce catastrophic forgetting, in which the model loses some of the rich generative knowledge from its original training.
Parameter-Efficient Fine-Tuning (PEFT): A more economical alternative trains only a small set of additional parameters while keeping the backbone of the pre-trained model frozen. LoRA [39], employed by methods such as StitchDiffusion [57] and DiffPano [59], typifies this class. The chief advantage is a drastic reduction in computational and memory requirements, making adaptation more accessible and largely preserving the original model’s powerful priors. The main limitation is that the frozen backbone can restrict the model’s ability to learn fundamentally new concepts, such as complex spherical geometry, making it potentially less expressive than a fully fine-tuned model.
Training-Free: The third strategy completely bypasses fine-tuning. Training-free methods like PanoFree [41] and SphereDiff [22] use pre-trained text-to-image models directly. They achieve panorama generation through careful orchestration of existing capabilities, such as iterative inpainting, view stitching, and specialized sampling algorithms. This approach requires no 360-degree panoramic training data and avoids costly training, fully leveraging the inherent power of the base models. The trade-offs include typically slower inference due to their multi-step, patch-based processes and a higher risk of global inconsistency, as there is no end-to-end training to enforce panoramic coherence.
IV-D Comparisons
To systematically evaluate strengths and weaknesses of representative methods, we conduct a benchmark on methods with publicly available inference code and model checkpoints. Our primary objective is to provide a fair and consistent comparison of the official releases, which requires running all models on the same hardware and evaluating them using a consistent set of metrics. We acknowledge that this approach introduces a selection bias, which means our quantitative comparison does not include several recent and important contributions that lack public code or model checkpoints. While this constraint limits the breadth of our comparison, it critically ensures the validity and fairness of the presented results. We encourage readers to consult the original papers for the reported performance of methods not included here. For Text-Only Generation, we compare Text2Light [56], Diffusion360 [58], StitchDiffusion [57], PanFusion [59], PAR [25], and SMGD [81]. For Text-Driven NFoV Outpainting, we compare PanoDiff [42], Diffusion360 [58], SpND [82] and DreamCube [83].
IV-D1 Qualitative Comparison
Fig. 6 presents a visual comparison of the six text-only generation methods. To facilitate the assessment of visual and semantic continuities at the image borders, we generate ”swapped” versions by exchanging the left and right halves of each generated image. Red dashed boxes are employed in these swapped images to highlight the boundary regions. Among the evaluated methods, Text2Light exhibits obvious seams in the highlighted areas of its swapped images, indicating a failure to maintain visual and semantic continuities at the borders. While Diffusion360 effectively mitigates such observable seams and preserves visual continuity, it occasionally demonstrates local semantic discontinuity at the borders, where content across image boundaries lacks consistent logical or meaningful coherence. In contrast, the other four methods achieve strong visual and semantic continuities at the borders of their synthesized images.
Fig. 7 displays the visual comparison for the four text-driven NFoV outpainting approaches. Similar to the text-only comparison, swapped image versions were created to scrutinize the visual and semantic continuities at the borders. PanoDiff, Diffusion360, and DreamCube demonstrate commendable visual and semantic continuities at the borders of the produced images, while SpND exhibits fair visual and semantic continuities.
@@ IV-D1 Qualitative Comparison
| SpND | 69.54 | 3.00 | 3.77 | 19.17 | 8.67 | 119.05 | 1.40 | 22 | 16.7 |
| DreamCube | 66.15 | 2.05 | 4.88 | 19.26 | 15.87 | 115.52 | 1.10 | 12 | 16.7 |
IV-D2 Quantitative Comparison
While the qualitative examples in Figs. 6 and 7 offer valuable visual insights, a quantitative evaluation is essential for a rigorous and objective comparison. To this end, we conduct a comprehensive comparison using metrics outlined in Sec. III-C.
@@ IV-D2 Quantitative Comparison
To ensure an unbiased evaluation of the generalizability of the methods, we employ an out-of-domain dataset, ODI-SR [1]. This dataset was specifically chosen for two primary reasons: (1) None of the evaluated models were trained on it, which guarantees a fair test of generalization to unseen data; and (2) its diverse composition of indoor and outdoor 360-degree panoramas provides a robust benchmark for evaluating performance across varied real-world scenarios. For generating text descriptions, we use BLIP2 [18] to create textual captions for the 360-degree panoramas included in the ODI-SR dataset. These generated text prompts serve as inputs for both Text-Only Generation and Text-Driven NFoV Outpainting tasks. To simulate NFoV images, we first project the equirectangular 360-degree panoramas from the ODI-SR dataset into a cubemap format and then extract the front face of each cubemap. The original 360-degree panoramas from ODI-SR are designated as real images and used as ground truth for the computation of evaluation metrics.
@@
The quantitative results obtained from this comparative evaluation, across the seven evaluation metrics, are presented in Tab. IV. To provide insights into the computational efficiency of each method, we also report the inference time and GPU memory required to generate a 1024512 360-degree panorama on a consistent machine equipped with an RTX A6000 GPU. Note that our results only represent the performance of the publicly released version of each method on a specific GPU (RTX A6000, 48 GB). The actual performance may be influenced by hardware differences and implementation-specific optimizations. These quantitative findings effectively delineate the strengths and weaknesses of the evaluated approaches, offering valuable guidance for future research in this domain.
V Emerging Directions
This section details two emerging directions that are closely related to text-driven 360-degree panoramic image generation: text-driven 360-degree 3D scene generation and text-driven 360-degree panoramic video generation, respectively.
V-A 360-Degree 3D Scene Generation
@@ V Emerging Directions
Recent advances in text-driven 360-degree panorama generation [57, 58, 59] have catalyzed innovative methods for reconstructing 360-degree 3D scenes from textual descriptions. 360-degree panoramic images inherently capture both global contexts and geometric constraints of a scene, making them an essential intermediate representation for 3D scene generation. Consequently, recent text-driven 360-degree 3D scene generation methods [27, 30, 31, 32, 33, 90, 91, 92] use 360-degree panorama generation to bridge the gap between text prompts and 360-degree 3D scene reconstruction.
@@
As depicted in Fig. 8, these methods typically use a two-stage process: (1) 360-Degree Panorama Generation: generating a 360-degree panorama from the input text prompt using a fine-tuned LDM [63], and (2) 3D Scene Reconstruction: inferring a 3D representation, typically with 3D Gaussian Splatting (3DGS) [26], from the generated panorama and corresponding multi-view perspective images. Tab. V provides a comparative summary of the methods using text-driven 360-degree panorama generation techniques introduced in Sec. IV.
@@
Within this framework, emerging methods are primarily differentiated by (a) their choice of 360-degree panorama generators and (b) their strategies for extracting and utilizing 3D information. For instance, FastScene [31] and HoloDreamer [30] both employ Diffusion360 [58] to generate the initial 360-degree panorama depicting a scene from a given text prompt. FastScene [31] then synthesizes multi-view panoramas of this scene for specific camera poses using Coarse View Synthesis and Progressive Novel View Inpainting. With these synthesized multi-view panoramas, FastScene introduces Multi-View Projection to get their perspective views. The point clouds derived from these views are then used as input for 3DGS to reconstruct the 3D scene. HoloDreamer [30] enhances the Diffusion360-generated panorama with two distinct ControlNet-based LDMs [37] and a super-resolution network to create a high-resolution, stylized output. Subsequently, HoloDreamer initializes 3D Gaussians using point clouds derived from a reverse equirectangular projection of the high-resolution panorama combined with its corresponding depth information. Finally, a two-stage 3DGS optimization process is developed to refine the scene rendering, resulting in the desired 3D scene reconstruction.
@@
Furthermore, certain methods deviate from the reliance on Diffusion360 [58]. DreamScene360 [33] uses StitchDiffusion [57] to generate multiple 360-degree panorama candidates and then employs a self-refinement process to select the optimal candidate for initializing panoramic 3D Gaussians with a 3D geometric field. To facilitate visual feature correspondences between different views and maintain geometric consistencies during the 3DGS optimization process, semantic and geometric regularizations are applied. In contrast, SceneDreamer360 [32] uses a fine-tuned PanFusion [59] generator, coupled with a super-resolution module from [58], to produce a high-resolution (6K) panorama aligned with the input text prompt. It then uses optimization-based viewpoint selection to extract multi-view images, which are subsequently used for improved point cloud initialization, ultimately leading to 3DGS-based scene reconstruction.
@@
Other methods explore alternative panorama generation techniques. LayerPano3D [27] begins by generating four orthogonal perspective views with a fine-tuned text-to-image model [21]. These initial views are then combined with text-guided inpainting [63], and further processed by using a fine-tuned Diffusion360 [58] model to outpaint the polar regions, resulting in a reference 360-degree panorama. To handle occlusions in complex scenes, LayerPano3D [27] decomposes the reference panorama into multiple depth-based layers and uses a fine-tuned inpainter [59] to complete unseen content at each layer. These inpainted, layered panoramas then provide supervision for panoramic 3D Gaussian scene optimization.
V-B 360-Degree Panoramic Video Generation
Analogous to the natural evolution from text-to-image generation [12, 13, 21, 63] to text-to-video (T2V) generation [98, 99, 100, 101], recent progress in text-driven 360-degree panoramic image synthesis has spurred research into the more challenging task: text-driven 360-degree panoramic video generation. Its representative methods [93, 94, 95, 96, 97] are summarized in Tab. VI.
Most works in this area rely on specialized training. 360DVD [93] pioneers this task by first constructing a tailored dataset, WEB360, consisting of 2,114 360-degree panoramic video-text pairs. Using this dataset, it trains an adapter to enable a pre-trained T2V models [99] to synthesize 360-degree panoramic videos from provided text prompts. Inspired by PanoDiff [42], 360DVD adopts a latent rotation mechanism in the inference process to maintain the boundary continuity of the synthesized results. Limitations in 360DVD, particularly the lack of detailed motion descriptions in its dataset, have prompted further refinements. PanoDiT [95] addresses this by curating a higher-quality subset named PHQ360 and replacing the U-Net architecture with a Diffusion Transformer (DiT) [102] and a motion LoRA for improved generation. Similarly, the rotation mechanism from PanoDiff is used in the post-processing phase of PanoDiT to ensure continuity. Recognizing the critical role of large-scale and high-quality data, PanoWan [97] establishes PanoVid, a dataset with over 13,000 video-text pairs. It introduces a latitude-aware sampling technique to mitigate ERP distortions and fined-tune a DiT-based T2V model [103] with LoRA for generation.
In a different vein, VideoPanda [96] seeks to better leverage priors from pre-trained T2V models. Trained on WEB360, it introduces multi-view attention layers to synthesize multiple perspective video outputs, which are then stitched together to form the final 360-degree panoramic video. This approach avoids direction in the equirectangular projection space.
Contrasting with these trained-based methods, several approaches have explored trained-free generation. DynamicScaler [94] designs an offset-shifting denoiser and a panoramic projection technique to synthesize a low-resolution 360-degree panoramic video, which then provides global motion guidance for refining a high-resolution version. However, its reliance on the ERP latent representation can lead to discontinuities near the poles. To address this, SphereDiff [22] introduces a spherical latent representation and extends Multidiffusion [55] to the constructed spherical space, achieving a more uniform distribution and improving quality at the poles. A common challenge to these training-free methods is that their patch-based synthesis mechanism can introduce global inconsistencies. Future work could focus on incorporating global guidance into these frameworks to mitigate this issue.
Datasets for 360-Degree Panoramic Video Generation. While the datasets discussed in Sec. III-B are foundational for static 360-degree panoramic image generation, the task of 360-degree panoramic video generation requires specialized datasets that include temporal information. These datasets are crucial for training and evaluating models capable of producing coherent and immersive 360-degree panoramic video content. Key datasets in this domain include:
360-1M [106] is a large-scale dataset consisting of 1,076,592 360-degree videos, collected from YouTube and distributed across 15 categories. As it was not originally created for text-driven generation, this dataset does not provide paired textual descriptions.
PanoVid [97] is a high-quality dataset of over 13,000 video clips curated specifically for text-driven 360-degree panoramic video generation. The videos in PanoVid are collected from multiple sources, including WEB360, YouTube360, and 360-1M. Qwen-2.5-VL [109] was adopted to produce rich textual descriptions for the video content.